OpenStudy (anonymous):

A quadratic equation x^2+2hx+4=x has two equal roots.Find the possible values of h.

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (sidsiddhartha):

for equal roots discriminant=0

OpenStudy (anonymous):

yes

OpenStudy (sidsiddhartha):

consider this equation ax^2 + bx + c = 0 so discriminat=b^-4ac=0

OpenStudy (sidsiddhartha):

b^2-4ac=0

OpenStudy (sidsiddhartha):

ok?

OpenStudy (anonymous):

k

OpenStudy (sidsiddhartha):

now in your equation b=2h and c=4 just plug them in

OpenStudy (michele_laino):

Please, try to calculate the discriminant \[\Delta \] of your equation, then discuss the cases \[\Delta = 0\]

OpenStudy (sidsiddhartha):

and a=1

OpenStudy (sidsiddhartha):

so-\[(2h)^2-4*1*4=0\]

OpenStudy (sidsiddhartha):

make sense?

OpenStudy (anonymous):